11 What is the impact on cash flow from operations in the current year based on the change in operating assets and liabilitie12 Whats the forecasted capital expenditure based on the information below? Net PP&E beginning of period: 15,000 Net PP&E en13 Whats the retained earnings end of period based on the information below? Retained earnings beginning of period: 7,500 Ne14 Whats the depreciation expense based on the information below? Depreciation (percent of sales): 4% Revenues: 60,000 Gross15 Whats the forecasted EBIT based on the information below? Revenues: 56,000 Cost of goods sold: 32,000 • SG&A: 8,500 Depre
Answer #1

Change in accounts receivable from previous year = 1825- 1725 = 100

Change in inventories from previous year = 1785- 1535 = 250

Change in accounts payable from previous year = 1475- 1325 = 150

Now a positive change in accounts receivable and inventory means a decrease in cash flow and a positive change in accounts payable means an increase in cash flows

So, change in cash flow =Change in accounts payable-(Change in accounts receivable+Change in inventories)

= 150 -(100+250)

= - 200 (correct option is 1st option)